# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v1

@@ IV-A2 Fetch Push
We also consider a continuous control task based on the Fetch mobile manipulator [24]. The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”). The state and action spaces are continuous ($\mathcal{S}\subset\mathbb{R}^{55},a\in\mathbb{R}^{4}$). Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
@@ IV-A3 Robot Navigation
We extended the CarGoal-v0 environment [25] to create a more complex adversarial navigation task. The ego robot (“e”) must reach a green goal (“g”) while avoiding a violet chaser (“c”) moving with a fixed speed ($v_{c}=7.5\times 10^{-4}$). Static hazards (“h”) and movable vase (“v”) are present but not penalized. The state and action dimensions are $s\in\mathbb{R}^{88}$ and $a\in\mathbb{R}^{2}$, respectively. Ten distance-based predicates were defined, involving euclidean distances between the robot, the chaser, the goal and the hazards: $\psi_{\mathrm{ec}}$, $\psi_{\mathrm{eg}}$, $\psi_{\mathrm{eh}}$, $\psi_{\mathrm{ev}}$, $\psi_{\mathrm{gc}}$, $\psi_{\mathrm{gh}}$, $\psi_{\mathrm{gv}}$, $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$, $\psi_{\mathrm{vc}}$.
To test robustness to noise, we included predicates irrelevant to task success (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) and evaluated whether TLNet could eliminate them in its explanations.
@@ IV-B Policy Optimization
TLNet is a general framework to generate explanations of robot policies, though, we focused on policies optimized by deep reinforcement learning algorithms in this work. We used Stable-Baselines3 [26] for training all policies. For the CtF and navigation environments, PPO [27] was used; for the Fetch push task, we used SAC [28] with Hindsight Experience Replay (HER) [29]. Hyperparameters were taken from prior work [16, 30, 25]. Blue agent policies in CtF achieved over 97% success, except the roomba case, which reached 77.5% success in 1,000 evaluation episodes. For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explanability metrics.
TABLE I: Baseline Comparison of Representative Generated Explanations.
Scenarios TLNet (Ours) Greedy Top-3 Top-5
@@
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
@@
$0.5\mathcal{G}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
$\mathcal{F}[0.73\neg\psi_{\mathrm{ra,bt}}$
@@
$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
CtF Capture 0
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}[(0.33\psi_{\mathrm{ba,rf}}$
$\mathcal{F}[0.67\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
@@
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
$0.5\mathcal{F}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
$0.5\mathcal{G}[(0.15\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})\wedge 0.06\psi_{\mathrm{ba,rf}}$
@@
$\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$
@@
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
$0.5\mathcal{F}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
$0.5\mathcal{G}[(0.29\psi_{\mathrm{ba,rf}}\vee 0.04\psi_{\mathrm{ra,bf}}\vee 0.07\neg\psi_{\mathrm{ra,df}})\wedge$
@@
$\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$
CtF Roomba
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.35\psi_{\mathrm{ba,rf}}\vee 0.65\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
@@
$\mathcal{F}[1.0\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.48\neg\psi_{\mathrm{ba,bt}}$
@@
$0.5\mathcal{F}[1.0\psi_{\mathrm{bt}}]\wedge$
$0.5\mathcal{G}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
$0.5\mathcal{F}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
$0.5\mathcal{G}[1.0\neg\psi_{\mathrm{bd}}]$
$\mathcal{G}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}[1.0\psi_{\mathrm{bt}}]$
Robot Navi. $0.5\mathcal{F}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$
TABLE II: Baseline Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics.
@@
Scenarios Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5
@@ IV-C Experimental Results
Table I presents representative explanations for each scenario, comparing our method with three baseline. Across all scenarios, TLNet consistently inferred both task ($\mathcal{F}$) and constraint ($\mathcal{G}$) clauses, while the top-$k$ methods often captured only one of the two. This structural incompleteness stems from the fact that top-$k$ pruning lacks guarantees on preserving clause-level composition, despite yielding concise formulas with high classification accuracy.
Our method produced well-structured explanations across environments. For instance, task clauses included $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”) in CtF, and $\psi_{\mathrm{bt}}$ (“block reaches the target”) and $\psi_{\mathrm{eg}}$ (“robot reaches the target” in Fetch and robot navigation. Constraint clauses correctly included safety conditions like $\neg\psi_{\mathrm{ra,bf}}$ (“the red agent does not capture the blue flag”) or $\neg\psi_{\mathrm{bd}}$ (“the block is not dropped”), with the exception of the navigation domain, where structural guarantees were again lacking in the baselines.
@@
As shown in Table II, our method outperformed all baselines in conciseness, consistency, and strictness, with only marginal differences in classification accuracy. On average, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness. The lower accuracy in the CtF Roomba scenario can be attributed to the suboptimality of the learned policy, which reduced the separability between positive and negative trajectories. These results support the qualitative findings above, validating the explainability of our method’s outputs.
Across CtF scenarios, TLNet consistently captured the task $\psi_{\mathrm{ba,rf}}$. The capture and capture 0 scenarios yielded identical constraint structures: $\mathcal{G}(0.30\,\psi_{\mathrm{ba,rf}}\vee 0.70\,\neg\psi_{\mathrm{ra,bf}})$ and $\mathcal{G}(0.31\,\psi_{\mathrm{ba,rf}}\vee 0.69\,\neg\psi_{\mathrm{ra,bf}})$, reflecting their similar optimized behaviors despite reward differences. In contrast, the fight scenario reflected the learned avoidance strategy through the constraint $\mathcal{G}(0.33\,\psi_{\mathrm{ba,rf}}\vee 0.67\,\neg\psi_{\mathrm{ba,ra}})$, emphasizing avoidance of battles.
The inferred explanation for Fetch Push was: “With importance 0.5, eventually the block reaches the target; with importance 0.5, globally, the grip contacts the block (0.41) or the block is on the target (0.59)”. While the task was correctly captured, the constraint clause was only partially valid due to the use of smoothed min/max activations. Since the grip reaches the block early in the episode, negative robustness signals were underemphasized. Additionally, the predicate $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”) was not inferred, likely due to minimal violations in both positive and random-walk trajectories. This illustrates a limitation of the binary classification formulation and the trajectory sampling strategy. Future work could address this by using adversarial or policy-guided sampling to generate more informative negative examples.
For the navigation task, TLNet inferred a clear and aligned explanation: “With importance 0.5, eventually the robot reaches the goal; with importance 0.5, globally, it avoids the chaser.” The method successfully ignored irrelevant predicates (e.g., $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$), demonstrating robustness against unnecessary input predicates and alignment with the tasks and constraints.
TABLE III: Ablation Study of CtF Capture Scenario.
Metrics Full No reg. No pruning No filter
Accuracy 0.951 0.954 0.776 0.947
Conciseness 0.408 0.396 0.147 0.289
Consistency 0.325 0.233 0.2 0.175
Strictness 0.117 0.117 0.063 0.106
Table III presents ablation results on the CtF Capture scenario. The full pipeline achieved the highest conciseness, consistency, and strictness. A variant without regularization slightly outperformed in accuracy, which is expected, as regularization biases optimization toward simpler explanations. However, the accuracy trade-off was minimal ($<0.01$), while consistency improved by 39.4% and conciseness by 3.0%. Removing the predicate filter significantly reduced conciseness (27.2%) and consistency (46.1%), highlighting its importance. The most impactful component was pruning, particularly in reducing explanation size—nearly tripling conciseness. Overall, the ablation study confirms that each step of our simplification process plays a vital role in producing interpretable, structurally sound explanations without compromising performance.

### v2

@@ IV-A2 Fetch Push
We also consider a continuous control task based on the Fetch mobile manipulator [27]. The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”). The state and action spaces are continuous ($s\in\mathbb{R}^{55},a\in\mathbb{R}^{4}$). Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
@@ IV-A3 Robot Navigation
Finally, we extended the CarGoal-v0 environment [28] to create a more complex adversarial navigation task. The ego robot (“e”) must reach a green goal (“g”) while avoiding a violet chaser (“c”) moving with a fixed speed ($v_{c}=7.5\times 10^{-4}$). Static hazards (“h”) and a movable vase (“v”) are present but not penalized. The state and action dimensions are $s\in\mathbb{R}^{88}$ and $a\in\mathbb{R}^{2}$, respectively. Ten distance-based predicates were defined, involving euclidean distances between the robot, the chaser, the goal, and the hazards: $\psi_{\mathrm{ec}}$, $\psi_{\mathrm{eg}}$, $\psi_{\mathrm{eh}}$, $\psi_{\mathrm{ev}}$, $\psi_{\mathrm{gc}}$, $\psi_{\mathrm{gh}}$, $\psi_{\mathrm{gv}}$, $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$, $\psi_{\mathrm{vc}}$.
We included predicates irrelevant to task success in each environment (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) to evaluate whether the considered methods could eliminate them from their explanations.
@@ IV-B Policy Optimization
TLNet is a general framework to generate explanations of robot policies, though, we focused on policies optimized by deep reinforcement learning algorithms in this work. We used Stable-Baselines3 [29] for training all policies. For the CtF and navigation environments, PPO [30] was used; for the Fetch Push task, we used SAC [31] with hindsight experience replay (HER) [32]. Hyperparameters were taken from prior work [19, 33, 28]. For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explainability metrics. The time horizon ($H$) for each environment is 100 for CtF, 200 for Fetch Push, and 400 for Robot Navigation.
TABLE I: Comparison of the Most Frequently Inferred Explanations.
Scenarios TLNet (Ours) NN-TLI STONE (top-3) STONE (top-5)
@@
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
$\mathcal{F}_{I}^{\bar{w}}[0.73\neg\psi_{\mathrm{ra,bt}}$
@@
$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
@@
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[(0.15\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})\wedge 0.06\psi_{\mathrm{ba,rf}}$
@@
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$
@@
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[(0.29\psi_{\mathrm{ba,rf}}\vee 0.04\psi_{\mathrm{ra,bf}}\vee 0.07\neg\psi_{\mathrm{ra,df}})\wedge$
@@
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$
CtF Mixture
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.44\psi_{\mathrm{ba,rf}}\vee 0.56\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
@@
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}_{I}^{\bar{w}}[0.48\neg\psi_{\mathrm{ba,bt}}$
@@
CtF Capture 0
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[(0.33\psi_{\mathrm{ba,rf}}$
$\mathcal{F}_{I}^{\bar{w}}[0.67\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
@@
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{bd}}]$
$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
Robot Navi. $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
TABLE II: Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics. The best results are bolded.
@@
STONE STONE STONE STONE
Scenarios Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5
@@ IV-C Experimental Results
IV-C1 Benefits and Limitations of TLNet
Table I presents the most frequent explanation obtained by each method for each scenario. Across all scenarios, TLNet consistently inferred both task and constraint clauses, while also producing concise and relatively easy-to-understand explanations. In comparison, NN-TLI typically yielded verbose explanations that are difficult to understand, while STONE often captured only one temporal clause, resulting in incomplete explanations.
The explanations produced by TLNet are also qualitatively reasonable (and often correct). For example, for all CtF scenarios, TLNet correctly inferred the task clause as $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”)—no other method correctly inferred this task for every scenario. The inferred CtF constraints are also syntactically reasonable. For example, in the Capture scenario, the blue agent (global) constraint includes “red agent not on the blue flag” because the capture red policy only targets the blue flag, which required the blue agent to learn to prevent the red agent from reaching the blue flag. Other scenarios, such as Fight and Patrol, did not include this predicate in the inferred constraint because those red policies do not target the blue flag. In comparison, the CtF constraints inferred by NN-TLI are generally too verbose to understand, while STONE did not infer any constraints.
For Fetch Push, TLNet also correctly inferred the task clause as $\psi_{\mathrm{bt}}$ (“block reaches the target”), which no other method correctly inferred. While TLNet inferred part of the expected constraint ($\psi_{\mathrm{gb}},\psi_{\mathrm{bt}}$), it missed an expected constraint $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”), likely due to insufficient occurrences of the grip touching the obstacle in the sampled data. Future research could mitigate this issue by introducing more intelligent sampling strategies for generating the negative dataset (e.g., adversarial or targeted trajectory sampling). NN-TLI also inferred part of the expected constraint, but also missed the $\neg\psi_{\mathrm{go}}$ constraint. Meanwhile, both variations of STONE inferred an incorrect constraint of $\psi_{\mathrm{bt}}$ (“block reaches the target”), since this is actually the task and not possible to always satisfy.
For Robot Navigation, all methods correctly inferred the task clause but TLNet was the only one to infer the correct constraint clause as $\neg\psi_{\mathrm{ec}}$ (“the robot does not hit the chaser”).
@@ IV-C1 Benefits and Limitations of TLNet
As shown in Table II, TLNet also outperformed all baselines in conciseness, consistency, and strictness, with marginal differences in classification accuracy. Across all scenarios, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness. These results support the qualitative findings above, validating the explainability of our method’s outputs and the use of these explainability metrics.
IV-C2 Benefit of Including Weights
We now demonstrate the benefit of including weights within inferred explanations (e.g., using wSTL rather than STL [25] or LTL [16, 17, 18]) by examining how inferred weights differ across CtF scenarios. Consider the Capture and Mixture explanations in Table I. Both capture and mixture red policies target the blue flag, but the capture policy is more aggressive in doing so. The blue agent, in response, learns slightly different policies: a more defensive policy in Capture, prioritizing the protection of its flag, and a more aggressive policy in Mixture, focusing more on capturing the red flag. This policy difference is reflected in the weights in the constraint clause, where the $\neg\psi_{\mathrm{ra,bf}}$ predicate receives a higher weight in the Capture explanation (0.70) than in the Mixture explanation (0.56), due to the Capture policy’s prioritization of flag protection. Conversely, the Mixture policy shows a higher weight for the $\psi_{\mathrm{ba,rf}}$ predicate (0.44) than the Capture policy (0.30), due to the Mixture policy’s prioritization of flag capture. Note that $\psi_{\mathrm{ba,rf}}$ appears in both the task and constraint clauses in these explanations. This duplication results from our design choice to fix equal weights (0.5) on the task and constraint clauses, which forces TLNet to express prioritization between the task and constraint by including relevant task predicates in the constraint clause. Although one could alternatively include the task and constraints weights within the optimized parameters, we found that doing so introduced training instability.
Now, consider the Capture and Capture 0 scenarios, which differ in the reward to kill the red agent (0.5 vs. 0). While one might naively expect this difference to produce different policy explanations, it actually should not change the optimal blue agent policies because, to win the game, the blue agent should first defend its flag (by killing the red agent), since the red agent uses the capture policy. We see that TLNet successfully captured this nuance by inferring constraints with nearly identical weights, reflecting its robustness to minor reward differences that still lead to similar policies. NN-TLI and STONE did not infer similar constraints across these scenarios. More broadly, this result also demonstrates the importance of inferring explanations from actual observed behaviors of RL agents rather than simply looking at their specified reward function.
TABLE III: Ablation Study of TLNet in the CtF Capture Scenario.
Metrics Full No Filter No Regularization No Pruning
Conciseness 0.408 0.289 0.396 0.147
Consistency 0.325 0.175 0.233 0.2
Strictness 0.117 0.106 0.117 0.063
Accuracy 0.951 0.947 0.954 0.776
IV-C3 Importance of TLNet Components
Finally, the ablation study results shown in Table III show that filtering, regularization, and pruning are all important to the performance of TLNet. Removing predicate filtering notably reduced conciseness (by 27.2%) and consistency (by 46.1%). Removing regularization decreased conciseness (by 2.94%) and significanbly reduced consistency (by 28.3%) with marginal accuracy improvement. Finally, removing the pruning step nearly tripled conciseness and halved strictness.